Question 1. Your sister is watching TV, the sound blasting at 75dB. You yell to her to turn down the volume, and she lowers the intensity level to 60dB, by what factor has the power dropped?

Question 2. A travelling wave on a string obeys the following equation: y=0.27sin(12x-500t)mm for x in meters and t in seconds. Find the equation for the transverse;
(a) velocity and
(b) acceleration of the particle of the string at x=20cm
(c) what is the displacement of this exactly 4s?
1
Expert's answer
2021-09-01T11:20:07-0400

1)


"75-60=10\lg x\\x=31.6"

2)


"v=0.135\cos{(2.4-500t)}\frac{m}{s}\\
a=-\sin{(2.4-500t)}\frac{m}{s^2}\\y=0.27\cos{(2.4-500(4))}=0.24\ mm"